OpenStudy (anonymous):

Find all the zeros of the polynomial (x + 1)(x - 4)(x2 - 7)

OpenStudy (anonymous):

-1, 4 -1, 4, ± (radical) 7 0 No zeros

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm f(x)= (x + 1)(x - 4)(x^2 - 7) }\) \(\large\color{blue}{ \rm x + 1=0~~~~~,~~~~~x - 4=0~~~~~~or~~~~~~x^2 - 7=0 }\)

OpenStudy (solomonzelman):

Solve for zeros in each `x-1` `x+4` and `x²-7`. (All answers will be real numbers, but not all rationals )

OpenStudy (anonymous):

thank you!